A 640 kg sample of water at 0°C is cooled to -6°C, and freezes in the process. How much heat is liberated? (For water Lf = 334 kJ/kg and Lv 2257 kJ/kg. The specific heat for ice is 2050 J/kg∙K.)

Group of answer choices

A)14,100 kJ

B)176,000 kJ

C)222,000 kJ

D)320,000 kJ
Answer #1

total heat = m * lf + m c delta_T

= 640 * 334 + 640 * 2.05 * 6

= 213760 + 7872

= 221632 kJ

= approx 222000 kJ ...

so answer is option C